OpenStudy (anonymous):

Rewrite in simplest radical form X ^ 5/6 ________ X ^ 1/6 Show each step of your process

OpenStudy (anonymous):

\[\frac{ \sqrt[2]{5} }{ \sqrt[6]{1} }\]

OpenStudy (welshfella):

when you divide these you subtract the exponents for example a^4 / a^2 = a^(4-2) = a^2

OpenStudy (welshfella):

aaafter dividing write as radicals

OpenStudy (anonymous):

@welshfella so the exponent would be 4/6 ?

OpenStudy (anonymous):

would the answer be X ^ 4/6 ?

OpenStudy (welshfella):

yes but in simpler from it is x^ 2/3

OpenStudy (anonymous):

so thats the answer?

OpenStudy (welshfella):